OpenStudy (anonymous):

the perimeter of a triangle is 57 in. The Lengths of the sides are represented by three consecutive odd integers. Find the sides of the triangle.

OpenStudy (primeralph):

A+B+C = 57 A = B+2 = C+4. C+4+C+2+C = 57. Find C, then A, B.

OpenStudy (anonymous):

wouldnt it be something like 1+(x+1)+(x+2)+(x+3)=57????

OpenStudy (primeralph):

No. Odd numbers increase by 2.

OpenStudy (anonymous):

so it would be x+x+2+x+4

OpenStudy (primeralph):

Yes. Remember to use x to find the remaining two sides.

myininaya (myininaya):

Like we have 3 numbers The 3 numbers are consecutive odd numbers For example the following are consecutive odd numbers: 1,3,5,7,9,11 Notice that this can be written as 1,1+2,1+2(2),1+2(3),1+2(4),1+2(5) Do you see a pattern: This pattern can be simplified to 1,1+2,1+4,1+6,1+8,1+10 What if we let n be 1, then we can write this as n,n+2,n+4,n+6,n+8,n+10 But we have only 3 consecutive numbers in your problem so we will only need the first 3 from the above pattern, Your first odd number will be n Your second odd number will be n+2 Your third odd number will be n+4 So we have the perimeters is just us adding the sides So we have n+(n+2)+(n+4)=57 Or you drop the parentheses and just say n+n+2+n+4=57

OpenStudy (anonymous):

so that would be 3n+6=57 then 57-6=51.. so then we have 3n=51 51/3= 17... so would that be the answer??

myininaya (myininaya):

That is one side. The other sides are n+2 and n+4.

OpenStudy (primeralph):

There are three answers.

OpenStudy (anonymous):

17, 19, and 21???

OpenStudy (primeralph):

Check by adding them up.

OpenStudy (anonymous):

so it would be 17+2 and 17 +4???

OpenStudy (anonymous):

for the other 2 sides???

OpenStudy (anonymous):

thank yuo guys! :)

OpenStudy (primeralph):

Maybe. Check by adding the results. If you get 57, you're correct. You're welcome.